OpenStudy (anonymous):

Isosceles trapezoid ABCD has diagonals AC = 10x + 7 and BD = 2x + 41. Find the value of x. A. x = 11 B. x = 6 C. x = 4.25 D. x = 2.83

OpenStudy (anonymous):

10x + 7 = 2x + 41 8x + 7 = 41 8x = 34 x = 17/4 or 4.25